# Tensor formalism for predicting synaptic connections with ensemble modeling or optimization

### v2

@@ VII Generalizations: Multiple synapses, sign constraints, and Predicting neuronal responses
VIII Discussion
@@
It is useful to illustrate how our abstract theory applies to biological problems with an example. Consider binocular integration in the zebrafish pretectum Naumann et al. (2016); Kubo et al. (2014), which originally motivated the development of our theory. The pretectum is a retinorecipient brain area Kramer et al. (2019); Robles et al. (2014) that integrates behaviorally relevant visual motion signals Yildizoglu et al. (2020); Wang et al. (2020) over space Zhang et al. (2022) and time Bahl and Engert (2020); Dragomir et al. (2020) to stabilize the fish’s eyes Kubo et al. (2014); Portugues et al. (2014); Wu et al. (2020) and its physical position in space Naumann et al. (2016); Yang et al. (2022). Functional imaging experiments have identified neuronal populations that respond with non-orthogonal patterns to monocular and binocular motion cues Naumann et al. (2016); Kubo et al. (2014); Wang et al. (2019). Previous models have hypothesized inter-hemispheric patterns of synaptic connectivity that could account for these responses Naumann et al. (2016); Kubo et al. (2014), and multiple groups are interested to test such models against connectomics data Hildebrand et al. (2017); Svara et al. (2022). Our theory could be applied to go beyond the particularities of individual models to find the synaptic connections that are most indispensable for generating the functional responses. For instance, we can use our formalism to compute $W$-critical values for every synapse, thereby identifying the synapses that are predicted with certainty for any maximally allowed weight norm. Since this bound is a priori unknown, we can rank how indispensable the synapses are according to their $W$-critical values Biswas and Fitzgerald (2022). Our formalism also allows us to model more realistic biological constraints by considering elliptic weight bounds that account for different costs for synapses within a pretectal hemisphere, across hemispheres, or from the retina to the pretectum. If conservative predictions from ensemble modeling are successful, then it would be interesting to consider whether hypothetical optimization criteria can account for more of the experimentally observed structure. For instance, how close is the pretectal implementation to the one that minimizes the weight norm?
@@
Our tensor-based analyses and resulting geometric algorithm relate interestingly to core concepts and algorithms from optimization theory. Most directly, the optimization problems we studied correspond to quadratic programming problems Boyd and Vandenberghe (2004); Wright (1999), and we’ve shown how various predictions from ensemble modeling can be cast as quadratic programming problems as well. Here we were able to provide an insightful geometric algorithm via our comprehensive understanding of the solution space geometry Biswas and Fitzgerald (2022). The latter was possible because we considered the ${\cal P}\leq{\cal N}$ case, a regime of biological interest, and it would be interesting to investigate how much of our framework generalizes to ${\cal P}>{\cal N}$. Standard quadratic programming algorithms Boyd and Vandenberghe (2004); Wright (1999) operate on the dual formulation of the optimization problem, taking advantage of the Karush-Kuhn-Tucker (KKT) conditions Kuhn and Tucker (1951); Karush (1939) to introduce multiple auxiliary variables. Our algorithm is able to bypass such constructions and obtain results that are more directly interpretable in terms of the correlations of pre- and post-synaptic neuronal activities. Our formalism thus provides a new approach to solving a rather popular class of optimization problems that have found wide applicability in science and technology McCarl et al. (1977).
@@
Generalizing our analysis from quadratic functions to other convex functions could be another fruitful direction, as convex optimization objectives appear in many scientific problems Boyd and Vandenberghe (2004). Firstly, our correspondence between $Q_{\mathtt{cr}}$ in an ensemble modeling approach with $Q_{\mathtt{\min}}$ obtained from an appropriate optimization generally holds true. Second, the insight that semi-constrained dimensions behave either as constrained or unconstrained pattern when minimizing an objective function carries over from quadratic functions to more general functions, as can be seen from the KKT conditions Kuhn and Tucker (1951); Karush (1939). Thus, we can reduce the the nonlinear optimization problem to an appropriate linear optimization problem. Since we have a complete parameterization of the solution space in terms of the $\eta$ coordinates Biswas and Fitzgerald (2022), it should be possible to use standard techniques from calculus to find minima of the convex function for the different possible equivalent linear problems. Note that a local minimum of a convex function is its unique global minimum. One could then in principle find the correct minimum using the brute force combinatorial approach described in Section VI. However, it would be better to leverage our geometric understanding of the solution space to generalize our geometric algorithm. Whether this is possible for arbitrary convex functions remains an open question.

### v3

@@ VII Generalizations: Multiple synapses, sign constraints, and Predicting neuronal responses
VIII Applications to optokinetic and optomotor stimulus responses in larval zebrafish
It is useful to illustrate how our abstract theory applies to biological problems with an example. Consider binocular integration in the zebrafish pretectum Naumann et al. (2016); Kubo et al. (2014), which originally motivated the development of our theory. The pretectum is a retinorecipient brain area Kramer et al. (2019); Robles et al. (2014) that integrates behaviorally relevant visual motion signals Yildizoglu et al. (2020); Wang et al. (2020) over space Zhang et al. (2022) and time Bahl and Engert (2020); Dragomir et al. (2020) to stabilize the fish’s eyes Kubo et al. (2014); Portugues et al. (2014); Wu et al. (2020) and its physical position in space Naumann et al. (2016); Yang et al. (2022). Functional imaging experiments have identified neuronal populations that respond with non-orthogonal patterns to monocular and binocular motion cues Naumann et al. (2016); Kubo et al. (2014); Wang et al. (2019). Previous models have hypothesized inter-hemispheric patterns of synaptic connectivity that could account for these responses Naumann et al. (2016); Kubo et al. (2014), and multiple groups are interested to test such models against connectomics data Hildebrand et al. (2017); Svara et al. (2022). Our theory could be applied to go beyond the particularities of individual models to find the synaptic connections that are most indispensable for generating the functional responses. In this context, we note that although our previous work Biswas and Fitzgerald (2022) determined the space of networks consistent with observed neuronal patterns, the framework and results presented in this paper make it possible to obtain optimization/ensemble-based predictions with correlated biological neuronal patterns.
To illustrate and test our formalism, here we focus on verifiable activity predictions that we can make by analyzing neuronal activity data in larval zebrafish in response to optokinetic  Kubo et al. (2014); Svara et al. (2022) and optomotor  Naumann et al. (2016) visual cues. Future work will make and test anatomy predictions using function-linked connectomics data  Svara et al. (2022). We will first discuss how we plan to model the neuronal circuit, which involves the retina and the pretectum, in the context of our framework. Next we will focus on the circuit responsible for optokinetic gaze stabilization in larval zebrafish. Finally, we’ll discuss how a very similar analysis can be applied to investigate the optomotor response, a behavior that enables a larval zebrafish to maintain its location by swimming against water currents.
VIII.1 Neural network model of binocular integration in the pretectum
Both Kubo et al. (2014) and Naumann et al. (2016), studying optokinetic and optomotor responses respectively, envisioned a recurrently connected network of pretectal neurons receiving four different input patterns from direction-selective retinal ganglion cells  Robles et al. (2014); Kramer et al. (2019). Both studies identified pretectal neurons that tended to respond similarly to each other to all the different visual stimuli in their experiments and accordingly assigned them to belong to the same neuronal population. To apply our formalism, we therefore decided to model the pretectal circuit as a recurrent network of ${\cal D}$ driven neuronal populations receiving feedforward connections from ${\cal I}=4$ input direction-selective retinal ganglion cell populations [^15] Each of these populations should be conceptualized as the distributed response of cells responding to one of the four patterns.. The optokinetic study identified ${\cal D}=11$ such functionally-defined neuronal populations, while the optomotor study found that the population response types were highly lateralized and identified ${\cal D}=2\times 8=16$ neuronal populations, 8 in each hemisphere [^16] It is possible that there is a different classification scheme of neuronal populations that is more appropriate for our formalism in linking structure and function, but we leave this investigation for future..
To summarize the network’s response properties, we constructed a ${\cal P}\times{\cal N}_{\mathtt{tot}}$ dimensional response matrix for each experiment, $z=\{z^{\mu}{}_{m},\ \mu=1,\dots,{\cal P},\mbox{ and }m=1,\dots,{\cal N}_{\mathtt{tot}}\}$, where ${\cal P}$ represents the number of stimulus conditions for the experiment, which was 8 in both cases, and ${\cal N}_{\mathtt{tot}}$ represents the sum total number of inputs that the pretectal population receives (we will often refer to “populations” as “neurons” for brevity). Apart from the ${\cal D}$ pretectal neurons and the ${\cal I}$ retinal inputs, we allowed ${\cal N}_{\mathtt{tot}}$ to also include a bias that can alter the threshold of the neurons to become active. Thus, ${\cal N}_{\mathtt{tot}}={\cal D}+{\cal I}+1$. Each pretectal population response was quantified as the average fluorescence measured during the stimulus presentation. Since fluorescence data were not available for the retinal inputs, we modeled each retinal population as an indicator function for whether the corresponding stimulus component was present. In other words, each population’s activity took some nonzero value whenever a corresponding monocular stimulus component was present in the binocular stimulus and was zero otherwise. We chose the response value to equal the average of all the positive responses of the pretectal populations over stimulus conditions [^17] Choosing other values for this constant is equivalent to altering the cost of connectivity between retina and pretectum relative to the cost of recurrent connections within the pretectum. Our formalism certainly allows us to consider such elliptic bounding/optimization functions. However, since our primary goal here is to present a proof-of-concept demonstration of the utility and applicability of our methods, we will stick to the above assignment of retinal input values as a way of approximately equalizing costs of different types of connections.. Finally, we modeled the bias term as a constant input that was present for all stimulus conditions, and we chose this constant to be the same as the value we chose for the retinal inputs [^18] It could also be interesting to explore the implications of changing this bias value in the future..
While comprehensive connectivity measurements that would allow us to test structural predictions based on ensemble modeling or optimization are not yet available, the general framework we have developed in this paper also allows us to predict the activity of neurons in response to a new input pattern. Therefore, we decided to test our formalism in the following way. We considered one pretectal neuron at a time (referred to as the target neuron) and treated its ${\cal N}$ input pathways from other pretectal populations, direct retinal inputs, and its bias as feedforward inputs to the target neuron. All the input-output response patterns for the target neuron were modeled as fixed points of neural network dynamics and can be read off from the $z$ matrix [^19] See Methods and Biswas and Fitzgerald (2022) for more details on how predictions can be made in a recurrent network by decomposing the network’s fixed point response patterns into ${\cal D}$ feedforward network response patterns, one for each target neuron.. We next used ${\cal P}-1=7$ input-output patterns to predict the response of the target neuron (on or off) to the held-out eighth pattern. For a given target neuron we repeated this procedure ${\cal P}=8$ times, where each time we predicted the response to a different held-out pattern.
More explicitly, we computed $W_{\mathtt{cr},\vec{x}^{\nu}}^{p}$, the weight norm below which all the incoming weight vectors onto the $p^{th}$ target neuron consistently predict an on or an off response to the $\nu^{th}$ input pattern when constrained to reproduce the $\mu\neq\nu$ response patterns. As prescribed in the previous section, it is the weight vector with the smallest norm that satisfies
$y^{\mu}=\Phi\left(\sum_{m=1}^{\cal N}x^{\mu}{}_{m}w_{p}{}^{m}\right)\ \forall\ \mu\neq\nu$ (119)
and
$\sum_{m=1}^{\cal N}x^{\nu}{}_{m}w_{p}{}^{m}=0\ ,$ (120)
where $y^{\mu}$ denotes the activity of the target neuron in stimulus condition $\mu$, $x^{\mu}{}_{m}$ denotes the activity of the input pathway $m$ in condition $\mu$, both $y^{\mu}$ and $x^{\mu}{}_{m}$ can be read off of $z$, and $\vec{w}_{p}$ is the vector of weights onto the target neuron. We also remind the readers that the predicted activity state (whether the target neuron responds or not to the $\nu^{th}$ stimulus) is given by the sign of the input drive from the minimum norm weights that can reproduce the target neuron’s responses to the other seven stimulus conditions,
$d^{\nu}\equiv\sum_{m=1}^{\cal N}x^{\nu}{}_{m}w_{\min}{}^{m}\ ,$ (121)
where we notationally neglect the dependence of $\vec{w}_{\min}$ on $p$ and $\nu$ for brevity. We would predict the target neuron to respond if $d^{\nu}>0$ and to not respond if $d^{\nu}\leq 0$. Having calculated the $W$-critical values for all the patterns, we ranked them in decreasing order for each target neuron. We hypothesized that patterns with higher ranks (higher $W$-criticals) will be consistent with observations.
We performed these calculations for every pretectal neuron and every (held-out) pattern. Our predictions for all the pretectal neurons, $m=1,\dots,{\cal D}$, and for all the stimulus conditions, $\nu=1,\dots,8$ for both the optokinetic and optomotor experiments, are depicted in Fig. 7. We also decided to compare these ensemble modeling based predictions with naive predictions that rank predictions based on the absolute values of the input drives without accounting for the full geometry of the solution space. These comparative predictions are also shown in Fig. 7. Note that the details of Fig. 7 will be explained panel-by-panel in subsections below, and here we merely mean to convey the overall logic of the analyses.
VIII.2 Including self coupling terms
Before discussing our results for optokinetic and optomotor activity data, there is an important biological issue related to self couplings that we wish to address. Since we are dealing with neuronal populations and not individual neurons, synapses between neurons within a given population would effectively manifest as self connections. However, we must ignore such self connections when computing held-out activity predictions, as including them would amount to using the response of the target neuron itself to “predict” its response to the held-out pattern. Nevertheless, we will now explain why these self-connections, even if they exist, do not affect the predictions ma when ignoring them.
Suppose $\vec{w}^{\prime}_{\min}$ is the weight-vector including self-connections that has the minimum norm among all weight vectors satisfying Eqns. 119 and 120, and recall that the norm of $\vec{w}^{\prime}_{\min}$ corresponds to $W_{\mathtt{cr},\vec{x}^{\nu}}^{p}$. Now, we can always decompose $\vec{w}^{\prime}_{\min}$ into a synapse vector without self-couplings, $\vec{u}$, and a self-synapse part, $u^{\cal N}$,
$\vec{w}^{\prime}_{\min}=\vec{u}+u^{\cal N}\widehat{e}_{\cal N}\Rightarrow||\vec{w}^{\prime}_{\min}||^{2}=||\vec{u}||^{2}+(u^{\cal N})^{2}\ ,$ (122)
where we’ve assigned the self-synapse to index ${\cal N}$ for notational convenience. Now for any given value of $u^{\cal N}$, there may be many different $\vec{u}$’s that give rise to weight vectors satisfying Eqns. 119 and 120, but the $\vec{u}$ defined by Eqn. 122 must be such that the resulting $\vec{w}^{\prime}$ has the smallest norm. Since the right hand side of Eqn.(122) is a sum of two positive terms, this means that, for the given value of $u^{\cal N}$, $||\vec{u}||$ must be minimized. We will therefore try to find this minimum norm $\vec{u}$ for a given value of $u^{\cal N}$.
For any pattern imposing an equality constraint [^20] This includes $\nu$ because of Eqn. 120., we note that $\vec{u}$ must satisfy
$y^{\mu}$ $=$ $\sum_{m=1}^{\cal N}x^{\mu}{}_{m}w_{p}{}^{m}=\sum_{m=1}^{{\cal N}-1}x^{\mu}{}_{m}u^{m}+y^{\mu}u^{\cal N}\ ,$ (123)
where we’ve noted that $x^{\mu}{}_{\cal N}=y^{\mu}$. This implies
$y^{\mu}$ $=$ $\sum_{m=1}^{{\cal N}-1}x^{\mu}{}_{m}\left({u^{m}\over 1-u^{\cal N}}\right)=\sum_{m=1}^{{\cal N}-1}x^{\mu}{}_{m}u^{\prime}{}^{m}\ ,$ (124)
where we have defined the rescaled input weights to the target neuron as
$u^{\prime}{}^{m}={u^{m}\over 1-u^{\cal N}}$ (125)
and assumed $u^{\cal N}<1$ to ensure that Eqn. 125 is well defined and avoid unstable amplification of neural activity [^21] Recall that $u^{\cal N}$ notates the amplitude of the target neuron’s self-synapse within a recurrent neural network..
For the semi-constrained patterns, we have $y^{\mu}=0$, which implies
$\sum_{m=1}^{{\cal N}}x^{\mu}{}_{m}w_{p}{}^{m}=\sum_{m=1}^{{\cal N}-1}x^{\mu}{}_{m}u^{m}\leq 0$ (126)
and
$\sum_{m=1}^{{\cal N}-1}x^{\mu}{}_{m}\left({u^{m}\over 1-u^{\cal N}}\right)=\sum_{m=1}^{{\cal N}-1}x^{\mu}{}_{m}u^{\prime}{}^{m}\ \leq 0\ ,$ (127)
where we’ve used $u^{\cal N}<1\Rightarrow 1-u^{\cal N}>0$ to ensure that the weight rescaling does not flip the inequality constraint.
Eqns. 124 and 127 together imply that $\vec{u}^{\prime}$ obeys the same nonlinear equation as the network weights without self-couplings (Eqns. 119 and 120). Moreover since, $||\vec{u}||\propto||\vec{u}^{\prime}||$ for given $u^{\cal N}$, the former is minimized when the latter is minimized, so that
$||\vec{u}||_{\min}=(1-u^{\cal N})||\vec{u}^{\prime}||_{\min}=(1-u^{\cal N})||\vec{w}_{\min}||\ ,$ (128)
where $\vec{w}_{\min}$ now denotes the minimum norm weight vector without self-synapses that satisfies all of the constraints. Thus for a given value of $u^{\cal N}$, the minimum norm with self-synapses is given by
$||\vec{w}^{\prime}_{\min}||^{2}=F(u^{\cal N})=(1-u^{\cal N})^{2}||\vec{w}_{\min}||^{2}+(u^{\cal N})^{2}\ .$ (129)
To find the value of $u^{\cal N}$ that minimizes this function we can equate
${dF\over du^{\cal N}}=0\ ,$ (130)
which after straightforward algebra yields
$u^{\cal N}={||\vec{w}_{\min}||^{2}\over 1+||\vec{w}_{\min}||^{2}}\ ,\mbox{ and}||\vec{w}^{\prime}_{\min}||^{2}={||\vec{w}_{\min}||^{2}\over 1+||\vec{w}_{\min}||^{2}}\ .$ (131)
We notice that the above relationship between $||\vec{w}^{\prime}_{\min}||^{2}$ and $||\vec{w}_{\min}||^{2}$ is monotonic, so the ranking of non-self-synapse $W$-critical values based on $||\vec{w}_{\min}||$ is identical to that based on $||\vec{w}^{\prime}_{\min}||$.
VIII.3 Optokinetic response predictions
Figure 7: Applying the theoretical framework to neuronal activity data. (A-I) Analysis of optokinetic response data from Svara et al. Svara et al. (2022). (A) Larval zebrafish were surrounded by LED monitors showing nasalward or temporalward motions to each eye. This experimentally elicited the optokinetic response. Adapted with permission from Kubo et al.  Kubo et al. (2014). (B) Eight monocular and binocular stimulus conditions were used in the experiment. Monocular: NL/- nasalward left eye, TL/- temporalward left eye, -/NR nasalward right eye, -/TR temporalward right eye; binocular: CW clockwise, CCW counterclockwise, FW forward, and BW backward. (C) Neuronal responses in pretectum were measured from the eleven most common response types identified by Kubo et al. Kubo et al. (2014). This included four monocular types (MoNL, MoTL, MoTR, MoNR) and seven binocular types (FEL, BEL, FER, BER, FSP, BSP, and FELR). Example calcium imaging traces shown at right. Each response type is identified with the barcode showing whether the neuron responded to each stimulus. For better visualization, here we have interchanged the pattern (here columns) and neuronal (here rows) indices as compared to the text. (D) Model architecture. All response types could receive feedforward inputs from left and right retinal neurons that respond to temporalward or nasalward motion (rtTL, rtNL, rtTR, rtNR) and recurrent inputs from other pretectrum response types. (E) We depict the $z$ matrix, which encodes the activity patterns of the pretectal response types, retinal inputs, and bias. (F) For each response type, we made activity predictions for a held-out stimulus condition based on minimum norm model that explains the other seven. Red indicates a positive input drive and neural response. Blue indicates a negative input drive that leads the neuron to not respond. (G) We plot $W^{p}_{cr,\vec{x}^{\nu}}$ values for each pairing of target neuron and held out stimulus. Larger values indicate more consistent responses across the solution space. (H) We ranked activity predictions based on the $W^{p}_{cr,\vec{x}^{\nu}}$ magnitudes (highest to lowest). The green boxes with the values on top indicate correct predictions and their ranking. The green boxes with the crosses show the first incorrect predictions. No predictions were assessed after the first incorrect one, indicated by white boxes. (I) Similar to H, we ranked activity predictions by the absolute values of the input drives until the first incorrect prediction. (J-N) Analysis of optomotor response data from Naumann et al. Naumann et al. (2016). (J) Moving monocular gratings were presented medially (towards the midline) and/or laterally (away from the midline) below larval zebrafish in their left and right visual fields (top). The stimulus set consisted of all monocular and binocular combinations of these components (bottom). Circle icons represent the motion shown to each eye. Black arrows represent the motion direction of each grating. Adapted with permission from Naumann et al. Naumann et al. (2016). (K) Model architecture. Retinal cells send projections contralaterally (rtLL - left eye, leftward motion; rtLR - left eye, rightward motion; rtRL - right eye, leftward motion; rtRR - right eye, rightward motion). In pretectum, neural response types were recurrently connected and lateralized to one side of the brain. Left-motion selective cells were usually localized to the left hemisphere, and right-motion selective cells were usually localized to the right hemisphere. oMIL and oMIR were lateralized oppositely, reflecting Naumann et al.’s hypothesis that these neurons relay lateral retinal signals from the contralateral eye across the pretectum. (L) Optomotor variant of (E). (M) Optomotor variant of (H). (N) Optomotor variant of (I).
In this subsection, we will discuss our analysis of the optokinetic neuronal response data in larval zebrafish and explain the results from our predictions. We used the Svara et al. dataset Svara et al. (2022) because its functional calcium imaging data are aligned to electron-microscopy-based structural data that could facilitate future anatomical tests of the theory. Larval zebrafish were stimulated in their left and/or right visual fields by visual gratings that moved in the nasal (i.e., tail-to-nose) or temporal (i.e., nose-to-tail) directions (Fig. 7A). Each of these four stimulus components (i.e., nasal motion in left visual field, temporal motion in left visual field, nasal motion in right visual field, temporal motion in right visual field) was presented as a monocular stimulus, and these components were also binocularly combined to mimic clockwise, counterclockwise, forward, and backward motions (Fig. 7B). The total number of stimulus conditions was thus ${\cal P}=8$. Earlier work from Kubo et al.  Kubo et al. (2014) identified ${\cal D}=11$ over-represented functional neuronal response types based on the responses of pretectal neurons to these stimuli (Fig. 7C), see Methods and  Kubo et al. (2014) for details. We therefore considered a retina-pretectum neural network model involving ${\cal D}=11$ pretectal populations and ${\cal I}=4$ monocular retinal input patterns (Fig. 7D).
Our ensemble modeling framework allowed us to make accurate predictions for neuronal responses to optokinetic stimuli. As explained previously and detailed in the Methods, we used the Svara et al. dataset Svara et al. (2022) to construct a response matrix, $z$, that included the stimulus responses of each pretectum response type, the four retinal inputs, and a stimulus-independent bias (Fig. 7E). We then considered each pairing of target neuron and held-out stimulus condition to compute the minimum norm input weights needed to explain the other ${\cal P}-1$ stimulus responses of the target neuron. Fig. 7F plots the corresponding input drive (Eqn. 121), and the sign of each matrix element determines the target neuron’s predicted activity state (active or inactive) in the held out stimulus condition based on either ensemble-modeling or weight-norm optimization. We then computed $W^{p}_{cr,\vec{x}^{\nu}}$ for each pairing of target neuron and held-out stimulus (Fig. 7G), with higher $W^{p}_{cr,\vec{x}^{\nu}}$ values indicating more consistency of the target neuron’s activity state across the ensemble of network models. Thus, for a given target neuron we ranked the stimulus conditions according to their $W^{p}_{cr,\vec{x}^{\nu}}$ values. We found that the top three ranked predictions were all correct across neuronal response types (Fig. 7H). For some neurons (e.g., MoNL), all of the correct predictions were rank-ordered to the front of the list.
This success of ensemble modeling is not guaranteed from individual models. For instance, out of the $11\times 8=88$ activity state predictions that could be made, 55 were accurately predicted by the minimum norm model. This is less than the chance level, as 63 correct predictions could be made by always predicting that the target neuron does not respond. The strongest magnitude predictions of the minimum norm model were also not reliable. When we ranked stimuli according to the absolute value of the input drive, we found that only the top-ranked predictions turned out to be consistently correct (Fig. 7I).
Achieving as many correct predictions as ensemble modeling without error is rather unlikely by random chance. The total number of ways one can rank-order the stimuli is ${\cal P}!$. For a given target neuron, suppose that $L^{2}$ optimization correctly predicts the activity state for ${\cal T}$ out of ${\cal P}=8$ patterns. If the rank-ordering of ensemble modeling made ${\cal E}$ correct predictions without making any mistakes, the number of ways this could happen is $(^{{\cal T}}_{{\cal E}}){\cal E}!({\cal P}-{\cal E})!$, where the first term represents the number of ways to select ${\cal E}$ correct predictions from the available set of ${\cal T}$, the second term represents the number of ways to order these correct predictions, and the third term represents the total number of ways to order the remaining ${\cal P}-{\cal E}$ correct and incorrect predictions. Thus, the probability that ensemble modeling would get at least ${\cal E}$ correct predictions by chance is given by
$P_{\mathtt{ensemble},i}={(^{{\cal T}}_{{\cal E}}){\cal E}!({\cal P}-{\cal E})!\over{\cal P}!}={{\cal T}!({\cal P}-{\cal E})!\over{\cal P}!({\cal T}-{\cal E})!}.$ (132)
The probability is tabulated as:
MoNL MoTL MoTR MoNR BSP FSP
0.018 0.071 0.071 0.071 0.018 0.018
FEL BEL BER FER FELR -
0.071 0.071 0.071 0.21 0.21 -
The median chance probability is thus 7.1%.
VIII.4 Optomotor response predictions
We similarly analyzed prectectal responses during the optomotor response, a visual behavior wherein larval zebrafish maintain a stable position by orienting to the optic flow and swimming against it Randlett et al. (2015); Naumann et al. (2016); Yang et al. (2022). We used an experimental dataset from Naumann et al. Naumann et al. (2016), which drove optomotor behavior by presenting ${\cal P}=8$ monocular and binocular combinations of leftward and rightward motion in the fish’s lower visual field (Fig. 7J) [^22] The static stimulus did not activate pretectal neurons and was shown between the other stimuli.. Naumann et al. used these stimuli to identify ${\cal D}=16$ over-represented functional response types that were grouped into 8 mirror-symmetric left-right pairs. Since pretectal responses to optomotor stimuli are highly lateralized with leftward (rightward) motion activating the left (right) pretectum  Naumann et al. (2016); Yildizoglu et al. (2020), we assumed that each pretectal response type appears only on one side of the brain. Moreover, we imposed the biological constraint that the retina projects contralaterally in zebrafish Robles et al. (2014). So in our model neural network (Fig. 7K), each pretectal population received inputs from ${\cal N}=18$ circuit elements, with the first 15 corresponding to the other pretectal populations, two corresponding to direction-selective retinal ganglion cells from the contralateral eye, and one bias term. The $z$ matrix collecting the responses of all ${\cal N}_{\mathtt{tot}}=21$ circuit elements is shown in Fig. 7L.
We again found that ensemble modeling made many accurate predictions. As with the optokinetic model, we rank-ordered our activity predictions either based on the magnitudes of $W$-critical or the absolute value of the input drive. The ensemble modeling predictions using $W$-critical perform remarkably well (Fig. 7M), predicting at least four activity states correctly before making an error. These predictions were better than the input-drive based predictions of the minimum norm model (Fig. 7N), which could only consistently make three correct predictions. The minimum norm model correctly predicting 94 out of 128 responses. Unlike the optokinetic model, this is better than chance levels, as only 78 correct predictions could be made by assuming that the neurons never respond. The chance probabilities of achieving the accuracy of ensemble modeling’s rank ordering was,
oBL oBR BL BR iBL iBR ioBL ioBR
0.036 0.036 0.036 0.107 0.036 0.036 0.25 0.25
iMmL iMmR MmL MmR oMlR oMlL sL sR
0.018 0.018 0.036 0.036 0.036 0.036 0.018 0.018
yielding a median chance percentage of only 3.6%.
IX Discussion
@@
We illustrated and tested our formalism by applying it to neuronal activity in the circuit responsible for binocular integration in the pretectum of larval zebrafish Kubo et al. (2014); Naumann et al. (2016); Svara et al. (2022). Our successful predictions demonstrate that our formalism is applicable to biological data and can provide useful information about how structure and function are linked in biological networks. Future work should explore whether even more accurate activity predictions can be made by considering alternate norms and neuron-type classifications. Future work should also apply our methods to make synapse-resolution anatomical predictions for pretectal connectivity that can be tested with emerging connectomics datasets Hildebrand et al. (2017); Svara et al. (2022). For instance, we can now use our formalism to compute $W$-critical values for every synapse, thereby identifying the synapses that are predicted with certainty for any maximally allowed weight norm. Since this bound is a priori unknown, we can rank how indispensable the synapses are according to their $W$-critical values Biswas and Fitzgerald (2022), just as we did while predicting activity-states. If conservative predictions from ensemble modeling are successful, then it would be interesting to consider whether hypothetical optimization criteria can account for more of the experimentally observed structure. For instance, we can assess how close the pretectal implementation is to the one that minimizes the Euclidean weight norm.
@@
Our tensor-based analyses and resulting geometric algorithm relate interestingly to core concepts and algorithms from optimization theory. Most directly, the optimization problems we studied correspond to quadratic programming problems Boyd and Vandenberghe (2004); Wright (1999), and we’ve shown how various predictions from ensemble modeling can be cast as quadratic programming problems as well. Here we were able to provide an insightful geometric algorithm via our comprehensive understanding of the solution space geometry Biswas and Fitzgerald (2022), as well as obtain useful analytical approximations. These were possible because we considered the ${\cal P}\leq{\cal N}$ case, a regime of biological interest, and it would be interesting to investigate how much of our framework generalizes to ${\cal P}>{\cal N}$. Standard quadratic programming algorithms Boyd and Vandenberghe (2004); Wright (1999) operate on the dual formulation of the optimization problem, taking advantage of the Karush-Kuhn-Tucker (KKT) conditions Kuhn and Tucker (1951); Karush (1939) to introduce multiple auxiliary variables. Our algorithm is able to bypass such constructions and obtain results that are more directly interpretable in terms of the correlations of pre- and post-synaptic neuronal activities. Our formalism thus provides a new approach to solving a rather popular class of optimization problems that have found wide applicability in science and technology McCarl et al. (1977). A comparison of the run times of the different algorithms and approximation methods revealed that our analytical solutions (e.g., the “complementary approximation”) were always fastest. Our geometric algorithm was faster than the standard MATLAB quadratic programming solver for small network sizes, but quadprog eventually overtook it. Our geometric algorithm remained a practical option for all network sizes considered. We emphasize that we value the geometric algorithm and analytical approximations more for the insights they provide than for their speeds.
@@
Generalizing our analysis from quadratic functions to other convex functions could be another fruitful direction, as convex optimization objectives appear in many scientific problems Boyd and Vandenberghe (2004). Firstly, our correspondence between $Q_{\mathtt{cr}}$ in an ensemble modeling approach with $Q_{\mathtt{\min}}$ obtained from an appropriate optimization generally holds true. Second, the insight that semi-constrained dimensions behave either as constrained or unconstrained pattern when minimizing an objective function carries over from quadratic functions to more general functions, as can be seen from the KKT conditions Kuhn and Tucker (1951); Karush (1939). Thus, we can reduce the nonlinear optimization problem to an appropriate linear optimization problem. Since we have a complete parameterization of the solution space in terms of the $\eta$ coordinates Biswas and Fitzgerald (2022), it should be possible to use standard techniques from calculus to find minima of the convex function for the different possible equivalent linear problems. Note that a local minimum of a convex function is its unique global minimum. One could then in principle find the correct minimum using the brute force combinatorial approach described in Section VI. However, it would be better to leverage our geometric understanding of the solution space to generalize our geometric algorithm. Whether this is possible for arbitrary convex functions remains an open question.